You want to obtain a sample to estimate a population proportion. Based on previous evidence, you believe the population proportion is approximately p∗=16%p∗=16%. You would like to be 99% confident that your estimate is within 5% of the true population proportion. How large of a sample size is required?
Answer #1

Given that, population proportion ( p* ) = 0.16

margin of error ( E ) = 0.05

A 99% confidence level has significance level (alpha) = 0.01 and critical value is, Zo/2 = 2.575

We want to find, the sample size ( n ),

» n = p* (1-p* ) (42 )2 = 0.16 * (1-0.16) * (UR)2 = 356.5 2.575 0.05

n approx 357

Therefore, required sample size is 357